When 250cm3 of saturated solution of copper (ii) tetraoxosulphate (vi) at 30°C was evaporated to dry lense 5g of the salt was obtained.What is the solubility of the salt at 30°C


1
Expert's answer
2021-05-14T02:35:00-0400

M(CuSO4) = 159.6 g/mol

n(CuSO4) "= \frac{5}{159.6}=0.0313 \;mol"

Proportion:

0.0313 mol – 250 mL

x – 1000 mL

"x = \frac{0.0313 \times 1000}{250}=0.125\; mol/L"

Answer: solubility 0.125 mol/L